OpenStudy (anonymous):

Work done by A in one day is haMrof the work done by B in one day. Work done by B is half of the work done by C in one day. If C alone can complete the work in 7 days, in how many days can A, B arid C together complete the work?

OpenStudy (anonymous):

let c = 1/7 work/day then b = 1/14 and a = 1/28 the total work/day is 1/7+1/14+1/28 = 1/4 it would take 4 days

OpenStudy (anonymous):

thank u